# Even-parity stability of hairy black holes in $U(1)$ gauge-invariant scalar-vector-tensor theories

### v1

@@ IV.2.2 Radial and angular Laplacian stability conditions
whose positivity is required for the Laplacian stability of the perturbation $V$. This forms our first angular Laplacian stability condition  [^6] The “3” appearing in the subscript of (91) was kept for historical reasons.. The remaining two propagation speeds associated with the gravitational and scalar DOFs generally couple with each other and can be expressed as
@@
For the absence of Laplacian instabilities associate with the perturbation $\psi$ and $\delta\phi$, we require that the squared propagation speeds given in Eq. (92) are real and non-negative. These conditions [with the assistance of the Vieta’s formula and Eq.(88)] are translated to give
@@ V.1 Stability analysis for the Model 1
Regarding the first Laplacian stability condition along the angular direction, we plug the full expressions of $v_{6}$, $v_{10}$, etc., into Eq. (91), we find that
@@
We proceed to the second angular Laplacian stability condition in which the story becomes a little sophisticated. To deal with this problem, we first convert $\mathcal{M}_{1}$ given in Eq. (93) into a dimensionless form by (i) solving Eqs. (115) for $A_{0}^{\prime}$  [^8] Notice that, this set of algebraic equations have 3 roots originally. However, only one of them is real so we keep it as our solution for $A_{0}^{\prime}$. and substituting it into $\mathcal{M}_{1}$; (ii) replacing the model parameter $\beta_{3}$ in $\mathcal{M}_{1}$ with the dimensionless ${\tilde{\beta}}_{3}$; (iii) changing the variable $r$ to the dimensionless one $\xi\equiv r_{h}/r$; (iv) taking advantage of the freedom for choosing unit system, setting $M_{\text{pl}}=r_{h}$ (cf., footnote #1). As a result, a dimensionless form of $\mathcal{M}_{1}$ is obtained as a function of $\{f,h,\xi,\kappa,{\tilde{\beta}}_{3}\}$. We postpone its lengthy explicit expression to Appendix B. Knowing the fact that $f,h\in[0,1]$, $\xi\in(0,1]$ since we have $r\in[r_{h},+\infty)$, $\kappa\in(0,\sqrt{2})$ and $|{\tilde{\beta}}_{3}|\lesssim{\cal O}(1)$Heisenberg et al. (2018), we can attempt to determine the range of $\mathcal{M}_{1}$ in the 5-dimension phase space spanned by $\{f,h,\xi,\kappa,{\tilde{\beta}}_{3}\}$. In practice, all of these could be fulfilled numerically on Mathematica. It turns out that the minimum of $\mathcal{M}_{1}$ in this phase space is indeed positive, which is about [^9] Strictly speaking, the true minimum should be bigger than this, since $f$ and $h$ are correlated so that certain regions of this phase space won’t be reached. $2.001$. This result guarantees the second angular Laplacian stability condition.
The third angular Laplacian stability condition (94) is even more complicated. To analyze $\mathcal{M}_{2}$, new strategies have to be selected. After some attempts, we know a posteriori that new constraints on ${\tilde{\beta}}_{3}$ have to be performed for this stability condition to hold. Now the mission is to determine the new constraint. For this goal and by referring the analysis to the second angular Laplacian stability condition in which the minimum of $\mathcal{M}_{1}$ is achieved near the $r=r_{h}$, as well as the stability conditions in odd-parity sector studied in Ref. Heisenberg et al. (2018) in which the previously mentioned constraint on ${\tilde{\beta}}_{3}$ is obtained near the $r=r_{h}$, we shall focus on the behavior of $\mathcal{M}_{2}$ near the horizon $r\to r_{h}$. As usual, a dimensionless form of $\mathcal{M}_{2}$ could be obtained by following the same procedure mentioned earlier. However, since we are focusing on the $r\to r_{h}$ limit, the phase space now reduces to 2 dimension, spanned by $\{\mu,{\tilde{\beta}}_{3}\}$. Now the $\mathcal{M}_{2}$ could be written as
@@
We plot it out in Fig. 1 of Appendix C. By looking at the contour of the allowed region for $\mathcal{M}_{2}$, we can read off the new upper bound for legal $|{\tilde{\beta}}_{3}|$, which is approximately given by $|{\tilde{\beta}}_{3}|\lesssim 0.441$. In addition, it’s worth mentioning here that, by adopting this new upper bound and the numerical method for analyzing $\mathcal{M}_{1}$ with Mathematica, the minimum of $\mathcal{M}_{2}$ in the allowed region of phase space $\{f,h,\xi,\kappa,{\tilde{\beta}}_{3}\}$ is indeed found to be positive. Thus, we conclude that, under the new constraint $|{\tilde{\beta}}_{3}|\lesssim 0.441$, the third angular Laplacian stability condition is guaranteed.
@@ V.2 Stability analysis for the Model 2
Moving to the second angular Laplacian stability condition, the story is even more sophisticated than that of Model 1 due to the presence of ${\tilde{\beta}}_{4}$. Because of this, the analysis of $\mathcal{M}_{1}$ will be divided into two parts. The first attempt is about the behavior around $r_{h}$. As usual, a dimensionless form of $\mathcal{M}_{1}$ is needed. The basic steps for obtaining that is the same as what we did in last subsection. The only difference is that the resultant $\mathcal{M}_{1}$ is now a function of six variables, i.e., $\{f,h,\xi,\kappa,{\tilde{\beta}}_{3},{\tilde{\beta}}_{4}\}$. On top of that, it became more accessible to expand $\mathcal{M}_{1}$ around the metric horizon. To the lowest order of $(r-r_{h})$, that leads to
@@
On using this expression, we can determine the minimum of $\mathcal{M}_{1}$ in the phase space of $\{\mu,{\tilde{\beta}_{3}},{\tilde{\beta}_{4}}\}$ with the built-in functions of Mathematica. Keeping in mind that $\mu\in(0,1)$, ${\tilde{\beta}_{3}}\in(-0.441,0.441)$ and ${\tilde{\beta}_{4}}\in(-0.125,0.251)$, it turns out that the minimum of $\mathcal{M}_{1}$ is a positive number, which supports the second angular Laplacian stability condition for Model 2. A similar treatment to $\mathcal{M}_{2}$ brings us a result in the form of Eq. (123). Because of its tedious expression, we put that in Appendix B. The result shows that the minimum of $\mathcal{M}_{2}$ in the phase space of $\{\mu,{\tilde{\beta}_{3}},{\tilde{\beta}_{4}}\}$ is also determined to be a positive number. Therefore, it seems like the currently known constraints are sufficient to support the third angular Laplacian stability condition.
@@
The solution for $f$ and $h$ are shown in the panel (a) of Fig. 3 in Appendix C. Since the upper limit of ${\tilde{\beta}_{3}}$ is updated to $\approx 0.441$ compared to Ref. Heisenberg et al. (2018), we adopted it with ${\tilde{\beta}_{4}}=0.1$ and $\mu=0.5$. We can find deviation between Model 2 and that of GR. With the solutions of $f$ and $h$ in hand, together with the previously mentioned algebraic expressions for $A_{0}^{\prime}$ and $\phi^{\prime}$, all the background quantities are known. Inserting them into ${\mathcal{M}}_{1}$ and ${\mathcal{M}}_{2}$, we further obtain their numerical values. This part of results are exhibited in the panel (b) of Fig. 3 in which ${\mathcal{M}}_{1}$ and ${\mathcal{M}}_{2}$ are plotted as functions of $r_{h}/r$ instead of $r/r_{h}$ to show a larger scope. Fig. 3 shows that, both these two quantities stay positive in the whole space outside the horizon. This result further supports that the second and third angular Laplacian stability conditions for Model 2 are satisfied111111 An ideal treatment is to test the positiveness of ${\mathcal{M}}_{1}$ and ${\mathcal{M}}_{2}$ using the Monte Carlo method and try to cover the whole allowed region in $\{\mu,{\tilde{\beta}_{3}},{\tilde{\beta}_{4}},\xi\}$ phase space. Nevertheless, as seen from, e.g., Eqs. (132)-(134), the complexity of our problem has already made the calculations quite time-consuming and it is preventing us from running a more complete analysis. In principle, much more computational resources are needed for executing such an analysis. .
@@ V.3 Stability analysis for the Model 3
We proceed to the second and third angular Laplacian stability conditions in which the story becomes astonishingly sophisticated, in comparing to that of Model 1. To conquer this problem, we shall first run the analysis around the $r=r_{h}$ where the algebraic equations in Eq. (126) can be solved for $A_{0}^{\prime}$ as $A_{0}^{\prime}(r=r_{h})=\sqrt{2\mu}=\kappa$ by choosing the unit system so that $M_{\text{pl}}=r_{h}$. On using this solution at the horizon, we start to expand $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ step by step around the horizon. After a sequence of time-consuming but straightforward manipulations, their expressions at the horizon converted into the dimensionless form were obtained. The one for $\mathcal{M}_{1}$ is given below
@@
while the one for $\mathcal{M}_{2}$ is shown in Appendix B due to its intricate structure. With the known constraints on ${\tilde{\beta}}_{3}$, ${\tilde{\beta}}_{4}$ and $\mu$ (so does $\kappa$), we can determine the minimum of $\left.\mathcal{M}_{1}\right|_{r\to r_{h}}$ and $\left.\mathcal{M}_{2}\right|_{r\to r_{h}}$ on the use of the built-in functions of Mathematica, which shows that both of them are positive so that the second as well as third angular Laplacian stability conditions are satisfied for the given parameter regions.
To further confirm these stability conditions, $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ will also be plotted out on $r\in[r_{h},+\infty)$ with chosen coupling constants. This could be fulfilled by choosing ${\mu}=0.5$, ${\tilde{\beta}}_{3}=0.441$ and ${\tilde{\beta}}_{4}=0.1$ as well as being given the corresponding numerical background solutions shown in the panel (a) of Fig. 4 in Appendix C. The fundamental steps and techniques of obtaining these numerical solutions are the same as what we used to draw the panel (a) of Fig. 3. The only difference is that, for Model 3 there is no analytic expression for $A_{0}^{\prime}$. Thus, $f$, $h$, and $A_{0}^{\prime}$ need to be numerically solved simultaneously. For this reason, all of these three variables are plotted in the panel (a) of Fig. 4. Again, to show a larger scope, all the quantities are plotted as functions of $\xi$. We note that $A_{0}^{\prime}$ is plotted in its dimensionless form $r_{h}^{-1}dA_{0}/d\xi$, which, as shown in the plot, equals $-\xi^{-2}A_{0}^{\prime}$. It’s worth mentioning here that, $A_{0}^{\prime}$ behaves as ${\cal O}(r^{-2})$ at the spatial infinity ($\xi\to 0$). That’s why the quantity $-\xi^{-2}A_{0}^{\prime}$ won’t blow up as $\xi\to 0$. The results of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are exhibited in the panel (b) of Fig. 4. It is clear that, both these two quantities stay positive at any distances outside the horizon. As a result, we conclude that this figure further supports the second and third angular Laplacian stability conditions in Model 3 inside the given parameter space.
@@
Angular first (91) Type II (117) Type II (LABEL:aLap1M2) Type III (128)
Laplacian second (93) Type III (132) Type IV  (123) & Fig.3 Type IV  (129) & Fig.4
third (94) Type III  (118) & Fig.1 Type IV (134) & Fig.3 Type IV (135) & Fig.4
@@ VI Conclusions
The situation becomes a little complicated when moving to the angular Laplacian conditions. First of all, we can easily confirm the first angular Laplacian condition in an analytic way for Model 1 and Model 2 as in Eqs. (117) and (LABEL:aLap1M2), respectively. However, for the second and third angular Laplacian conditions of Model 1, semi-analytic as well as numerical techniques become necessary. It is found that these stability conditions are guaranteed by applying the constraints to the coupling parameters obtained from the odd-parity sector in Ref. Heisenberg et al. (2018) together with the additional constraints indicated in Fig. 1 (see Appendix C), namely, ${\tilde{\beta}_{3}}\in(-0.441,0.441)$ and ${\tilde{\beta}_{4}}\in(-0.125,0.251)$. With these constraints, the first angular Laplacian condition for Model 3 can also be confirmed [cf., (128)]. Finally, we notice that the second and third angular Laplacian conditions for Model 2 and Model 3 are quite difficult to investigate (especially for the Model 3). To conquer this difficulty, by mimicking Ref. Heisenberg et al. (2018), specific values for the coupling parameters are chosen for these two models. As a result, we are able to plot out the discriminants [cf., (93) and (94)] for each case. These two stability conditions are then confirmed in Figs. 3 and 4 (cf., Appendix C). The main formulas and analysis about these stability conditions are summarized for the three models in Table 1.

### v2

@@ IV.2.2 Radial and angular Laplacian stability conditions
whose positivity is required for the Laplacian stability of the perturbation $V$. This forms our third angular Laplacian stability condition, as indicated in the subscript [^6] This is referred as the third angular Laplacian stability condition since it belongs to the vector DOF, which appears as the third one in (64). For a more convenient way of narration, this one is discussed before the first and second angular Laplacian stability conditions.. The remaining two propagation speeds associated with the gravitational and scalar DOFs generally couple with each other and can be expressed as
@@
For the absence of Laplacian instabilities associate with the perturbation $\psi$ and $\delta\phi$, we require that the squared propagation speeds given in Eq. (92) are real and non-negative. These conditions [with the assistance of the Vieta’s formula and Eq.(88)], with the nearest two below referred as the first and second angular Laplacian stability condition, respectively, are translated to give
@@ V.1 Stability analysis for the Model 1
Regarding the third Laplacian stability condition along the angular direction, we plug the full expressions of $v_{6}$, $v_{10}$, etc., into Eq. (91), we find that
@@
We proceed to the first angular Laplacian stability condition in which the story becomes a little sophisticated. To deal with this problem, we first convert $\mathcal{M}_{1}$ given in Eq. (93) into a dimensionless form by (i) solving Eqs. (115) for $A_{0}^{\prime}$  [^8] Notice that, this set of algebraic equations have 3 roots originally. However, only one of them is real so we keep it as our solution for $A_{0}^{\prime}$. and substituting it into $\mathcal{M}_{1}$; (ii) replacing the model parameter $\beta_{3}$ in $\mathcal{M}_{1}$ with the dimensionless ${\tilde{\beta}}_{3}$; (iii) changing the variable $r$ to the dimensionless one $\xi\equiv r_{h}/r$; (iv) taking advantage of the freedom for choosing unit system, setting $M_{\text{pl}}=r_{h}$ (cf., footnote #1). As a result, a dimensionless form of $\mathcal{M}_{1}$ is obtained as a function of $\{f,h,\xi,\kappa,{\tilde{\beta}}_{3}\}$. We postpone its lengthy explicit expression to Appendix B. Knowing the fact that $f,h\in[0,1]$, $\xi\in(0,1]$ since we have $r\in[r_{h},+\infty)$, $\kappa\in(0,\sqrt{2})$ and $|{\tilde{\beta}}_{3}|\lesssim{\cal O}(1)$Heisenberg et al. (2018), we can attempt to determine the range of $\mathcal{M}_{1}$ in the 5-dimension phase space spanned by $\{f,h,\xi,\kappa,{\tilde{\beta}}_{3}\}$. In practice, all of these could be fulfilled numerically on Mathematica. It turns out that the minimum of $\mathcal{M}_{1}$ in this phase space is indeed positive, which is about [^9] Strictly speaking, the true minimum should be bigger than this, since $f$ and $h$ are correlated so that certain regions of this phase space won’t be reached. $2.001$. This result guarantees this angular Laplacian stability condition.
The second angular Laplacian stability condition (94) is even more complicated. To analyze $\mathcal{M}_{2}$, new strategies have to be selected. After some attempts, we know a posteriori that new constraints on ${\tilde{\beta}}_{3}$ have to be performed for this stability condition to hold. Now the mission is to determine the new constraint. For this goal and by referring the analysis to the first angular Laplacian stability condition in which the minimum of $\mathcal{M}_{1}$ is achieved near the $r=r_{h}$, as well as the stability conditions in odd-parity sector studied in Ref. Heisenberg et al. (2018) in which the previously mentioned constraint on ${\tilde{\beta}}_{3}$ is obtained near the $r=r_{h}$, we shall focus on the behavior of $\mathcal{M}_{2}$ near the horizon $r\to r_{h}$. As usual, a dimensionless form of $\mathcal{M}_{2}$ could be obtained by following the same procedure mentioned earlier. However, since we are focusing on the $r\to r_{h}$ limit, the phase space now reduces to 2 dimension, spanned by $\{\mu,{\tilde{\beta}}_{3}\}$. Now the $\mathcal{M}_{2}$ could be written as
@@
We plot it out in Fig. 1 of Appendix C. By looking at the contour of the allowed region for $\mathcal{M}_{2}$, we can read off the new upper bound for legal $|{\tilde{\beta}}_{3}|$, which is approximately given by $|{\tilde{\beta}}_{3}|\lesssim 0.441$. In addition, it’s worth mentioning here that, by adopting this new upper bound and the numerical method for analyzing $\mathcal{M}_{1}$ with Mathematica, the minimum of $\mathcal{M}_{2}$ in the allowed region of phase space $\{f,h,\xi,\kappa,{\tilde{\beta}}_{3}\}$ is indeed found to be positive. Thus, we conclude that, under the new constraint $|{\tilde{\beta}}_{3}|\lesssim 0.441$, the second angular Laplacian stability condition is guaranteed.
@@ V.2 Stability analysis for the Model 2
Moving to the first angular Laplacian stability condition, the story is even more sophisticated than that of Model 1 due to the presence of ${\tilde{\beta}}_{4}$. Because of this, the analysis of $\mathcal{M}_{1}$ will be divided into two parts. The first attempt is about the behavior around $r_{h}$. As usual, a dimensionless form of $\mathcal{M}_{1}$ is needed. The basic steps for obtaining that is the same as what we did in last subsection. The only difference is that the resultant $\mathcal{M}_{1}$ is now a function of six variables, i.e., $\{f,h,\xi,\kappa,{\tilde{\beta}}_{3},{\tilde{\beta}}_{4}\}$. On top of that, it became more accessible to expand $\mathcal{M}_{1}$ around the metric horizon. To the lowest order of $(r-r_{h})$, that leads to
@@
On using this expression, we can determine the minimum of $\mathcal{M}_{1}$ in the phase space of $\{\mu,{\tilde{\beta}_{3}},{\tilde{\beta}_{4}}\}$ with the built-in functions of Mathematica. Keeping in mind that $\mu\in(0,1)$, ${\tilde{\beta}_{3}}\in(-0.441,0.441)$ and ${\tilde{\beta}_{4}}\in(-0.125,0.251)$, it turns out that the minimum of $\mathcal{M}_{1}$ is a positive number, which supports this angular Laplacian stability condition for Model 2. A similar treatment to $\mathcal{M}_{2}$ brings us a result in the form of Eq. (123). Because of its tedious expression, we put that in Appendix B. The result shows that the minimum of $\mathcal{M}_{2}$ in the phase space of $\{\mu,{\tilde{\beta}_{3}},{\tilde{\beta}_{4}}\}$ is also determined to be a positive number. Therefore, it seems like the currently known constraints are sufficient to support this angular Laplacian stability condition.
@@
The solution for $f$ and $h$ are shown in the panel (a) of Fig. 3 in Appendix C. Since the upper limit of ${\tilde{\beta}_{3}}$ is updated to $\approx 0.441$ compared to Ref. Heisenberg et al. (2018), we adopted it with ${\tilde{\beta}_{4}}=0.1$ and $\mu=0.5$. We can find deviation between Model 2 and that of GR. With the solutions of $f$ and $h$ in hand, together with the previously mentioned algebraic expressions for $A_{0}^{\prime}$ and $\phi^{\prime}$, all the background quantities are known. Inserting them into ${\mathcal{M}}_{1}$ and ${\mathcal{M}}_{2}$, we further obtain their numerical values. This part of results are exhibited in the panel (b) of Fig. 3 in which ${\mathcal{M}}_{1}$ and ${\mathcal{M}}_{2}$ are plotted as functions of $r_{h}/r$ instead of $r/r_{h}$ to show a larger scope. Fig. 3 shows that, both these two quantities stay positive in the whole space outside the horizon. This result further supports that the first and second angular Laplacian stability conditions for Model 2111111 An ideal treatment is to test the positiveness of ${\mathcal{M}}_{1}$ and ${\mathcal{M}}_{2}$ using the Monte Carlo method and try to cover the whole allowed region in $\{\mu,{\tilde{\beta}_{3}},{\tilde{\beta}_{4}},\xi\}$ phase space. Nevertheless, as seen from, e.g., Eqs. (132)-(134), the complexity of our problem has already made the calculations quite time-consuming and it is preventing us from running a more complete analysis. In principle, much more computational resources are needed for executing such an analysis. .
@@ V.3 Stability analysis for the Model 3
We proceed to the first and second angular Laplacian stability conditions in which the story becomes astonishingly sophisticated, in comparing to that of Model 1. To conquer this problem, we shall first run the analysis around the $r=r_{h}$ where the algebraic equations in Eq. (126) can be solved for $A_{0}^{\prime}$ as $A_{0}^{\prime}(r=r_{h})=\sqrt{2\mu}=\kappa$ by choosing the unit system so that $M_{\text{pl}}=r_{h}$. On using this solution at the horizon, we start to expand $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ step by step around the horizon. After a sequence of time-consuming but straightforward manipulations, their expressions at the horizon converted into the dimensionless form were obtained. The one for $\mathcal{M}_{1}$ is given below
@@
while the one for $\mathcal{M}_{2}$ is shown in Appendix B due to its intricate structure. With the known constraints on ${\tilde{\beta}}_{3}$, ${\tilde{\beta}}_{4}$ and $\mu$ (so does $\kappa$), we can determine the minimum of $\left.\mathcal{M}_{1}\right|_{r\to r_{h}}$ and $\left.\mathcal{M}_{2}\right|_{r\to r_{h}}$ on the use of the built-in functions of Mathematica, which shows that both of them are positive so that the first as well as second angular Laplacian stability conditions are satisfied for the given parameter regions.
To further confirm these stability conditions, $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ will also be plotted out on $r\in[r_{h},+\infty)$ with chosen coupling constants. This could be fulfilled by choosing ${\mu}=0.5$, ${\tilde{\beta}}_{3}=0.441$ and ${\tilde{\beta}}_{4}=0.1$ as well as being given the corresponding numerical background solutions shown in the panel (a) of Fig. 4 in Appendix C. The fundamental steps and techniques of obtaining these numerical solutions are the same as what we used to draw the panel (a) of Fig. 3. The only difference is that, for Model 3 there is no analytic expression for $A_{0}^{\prime}$. Thus, $f$, $h$, and $A_{0}^{\prime}$ need to be numerically solved simultaneously. For this reason, all of these three variables are plotted in the panel (a) of Fig. 4. Again, to show a larger scope, all the quantities are plotted as functions of $\xi$. We note that $A_{0}^{\prime}$ is plotted in its dimensionless form $r_{h}^{-1}dA_{0}/d\xi$, which, as shown in the plot, equals $-\xi^{-2}A_{0}^{\prime}$. It’s worth mentioning here that, $A_{0}^{\prime}$ behaves as ${\cal O}(r^{-2})$ at the spatial infinity ($\xi\to 0$). That’s why the quantity $-\xi^{-2}A_{0}^{\prime}$ won’t blow up as $\xi\to 0$. The results of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ are exhibited in the panel (b) of Fig. 4. It is clear that, both these two quantities stay positive at any distances outside the horizon. As a result, we conclude that this figure further supports the first and second angular Laplacian stability conditions in Model 3 inside the given parameter space.
@@
Angular third (91) Type II (117) Type II (LABEL:aLap1M2) Type III (128)
Laplacian first (93) Type III (132) Type IV  (123) & Fig.3 Type IV  (129) & Fig.4
second (94) Type III  (118) & Fig.1 Type IV (134) & Fig.3 Type IV (135) & Fig.4
@@ VI Conclusions
The situation becomes a little complicated when moving to the angular Laplacian conditions. First of all, we can easily confirm the third angular Laplacian condition in an analytic way for Model 1 and Model 2 as in Eqs. (117) and (LABEL:aLap1M2), respectively. However, for the first and second angular Laplacian conditions of Model 1, semi-analytic as well as numerical techniques become necessary. It is found that these stability conditions are guaranteed by applying the constraints to the coupling parameters obtained from the odd-parity sector in Ref. Heisenberg et al. (2018) together with the additional constraints indicated in Fig. 1 (see Appendix C), namely, ${\tilde{\beta}_{3}}\in(-0.441,0.441)$ and ${\tilde{\beta}_{4}}\in(-0.125,0.251)$. With these constraints, the third angular Laplacian condition for Model 3 can also be confirmed [cf., (128)]. Finally, we notice that the first and second angular Laplacian conditions for Model 2 and Model 3 are quite difficult to investigate (especially for the Model 3). To conquer this difficulty, by mimicking Ref. Heisenberg et al. (2018), specific values for the coupling parameters are chosen for these two models. As a result, we are able to plot out the discriminants [cf., (93) and (94)] for each case. These two stability conditions are then confirmed in Figs. 3 and 4 (cf., Appendix C). The main formulas and analysis about these stability conditions are summarized for the three models in Table 1.